one serving (1/2 cup) cooked peas contains 45 more calories than one serving cooked carrots and 50 m0re calories than one serving of cooked green beans. if one serving of carrots and three servings of green beans contain the same number of calories as one serving of peas, how many calories are there in one serving of peas.
Answer by mananth(16946)

one serving (1/2 cup) cooked peas contains 45 more calories than one serving cooked carrots and 50 m0re calories than one serving of cooked green beans. if one serving of carrots and three servings of green beans contain the same number of calories as one serving of peas, how many calories are there in one serving of peas.
1 serving beans x calories
1 serving carrots = y cal
1 serving peas = z cal
..
z=y+45...............1
z= x+50...............2
z=y+3x................3
..
multiply 1 by -1 andd to 3
-z=-y-45
sdd
z-z=-y+y-45+3x
3x=45
/3
x= 15 calories ---beans
..
z= x+50 =15+50
=65 cal--- peas
z=y+3x
65=y+45
y=65-45
y=20 cal ---carrots
..
m.ananth@hotmail.ca